OpenStudy (anonymous):

Two boxes are shaped like cubes. Each edge of the smaller box is 1/2 the length of the corresponding edge of the larger box. The larger box will hold. How many times as much sand as the larger box ? Plz help !!

Vocaloid (vocaloid):

let L = length of the edge for the large box let (1/2)L = length of the edge for the small box volume = (length of edge)^3 large box volume/small box volume = ?

OpenStudy (anonymous):

@Vocaloid should I do 1/2 times 3

Vocaloid (vocaloid):

^ means exponent so x^3 means x*x*x

OpenStudy (anonymous):

@Vocalaid I'm confused now

OpenStudy (anonymous):

@vocaloid

Vocaloid (vocaloid):

let's just find the volume of the large box first, ok? edge of large box = L volume = edge^3 = (edge)*(edge)*(edge) volume of large box = ?

OpenStudy (anonymous):

V= 27 ?

Vocaloid (vocaloid):

edge = L (edge)*(edge)*(edge) = edge^3 replace "edge" with "L"

OpenStudy (anonymous):

Kk

OpenStudy (anonymous):

@Vocaloid I still dont understand

Vocaloid (vocaloid):

edge = L volume = (edge)^3 to find volume, replace the word "edge" with "L" volume of large box = L^3 now do the same thing with the small box

OpenStudy (anonymous):

What does ^3 mean ???

OpenStudy (anonymous):

Its an exponent ?

Vocaloid (vocaloid):

yes...

Vocaloid (vocaloid):

x^3 = x*x*x

Vocaloid (vocaloid):

L^3 = L*L*L, etc....

OpenStudy (anonymous):

Kk I got it now . whats the next step ?

Vocaloid (vocaloid):

find the volume of the small box edge = (1/2)L volume = edge^3 = ?

OpenStudy (anonymous):

Its 27 ? Since 3×3×3

Vocaloid (vocaloid):

no, I don't know where you're getting 3 from... edge = (1/2)L volume = edge^3 = (1/2)L*(1/2)L*(1/2)L = ?

OpenStudy (anonymous):

Ohhhhhhhh sorry I kinda misunderstood what you meant I understand Now

OpenStudy (anonymous):

Ohhhhhhhh sorry I kinda misunderstood what you meant I understand Now

OpenStudy (anonymous):

@Vocaloid I got 1/8

Vocaloid (vocaloid):

you forgot the L L^3 = ?

OpenStudy (anonymous):

Yeah its equal to 1/8

Vocaloid (vocaloid):

(1/2)L*(1/2)*L*(1/2)L = (1/8)*__^3 fill in the blank you forgot to include the L

OpenStudy (anonymous):

1/8L

Vocaloid (vocaloid):

still missing the exponent

OpenStudy (anonymous):

V = 1/8³L ??

OpenStudy (anonymous):

V = 1/8³L ??